OpenStudy (anonymous):

if the exterior angle of a regular polygon equals 15 degrees how many sides does it have?

OpenStudy (anonymous):

If N is number of sides - Then exterior angle = 360/N But exterior angle = 15 15 = 360/N N = 360/15 = 24 sides.

OpenStudy (anonymous):

thanks man really confused with these questions